# Effective ML Model Versioning in Edge Networks

### v1

@@ VI Conclusion
In this paper, we addressed the problem of automating ML model versioning in constrained edge networks, considering different kind of updates to ML models. We compared a RL based approach to other update strategies, which today are common practice. The simulation results demonstrate that the quality of RL-based decision depends on the traffic load. For less load, RL achieves comparable results to the always update strategy. At the same time, it outperforms the random update in all measurements. Perhaps most importantly, our study showed important limitations and thus directions for further research. We assumed that subversion updates always improve (linearly or exponentially), which is practically not always the case. Also we did not take the stability as a values for the optimization problem into consideration yet, which also is premium in real world systems. Also many more, and larger topologies need to be studied, along with the proper optimizations of the solution. Finally, we did not take into account any downgrading of a version. These and many other interesting findings are subject of future research.
Acknowledgment
This work was partially supported by the EU project MANOLO under GA no. 101135782.

### v2

@@ VI Conclusion
In this paper, we addressed the problem of effective automation of ML model versioning in constrained edge networks. We compared a RL based approach to other update common strategies. The Update Agent, when implemented with reinfocement learning, showed a notable and practically relevant trend. A replica was more often updated in instances where multiple subversion updates were available, rather than deploying the most recent replica for each minor update. The simulation results demonstrate that the quality of RL-based decision depends on the server load. For less load, RL achieves comparable results to the always update strategy for ML model specific parameters, while achieving better performance. Perhaps most importantly, our study showed important limitations and thus directions for further research. We assumed that subversion updates always improve (linearly or exponentially), which is practically not always the case. We did not take the stability as a values for the optimization problem into consideration yet, which also is premium in real world systems. Also, we did not take into account any downgrading of a version. These and many other interesting findings are subject of future research.